OpenStudy (anonymous):

Please help solve the Initial Value Problem: x^2(dy/dx)=(4x^2-x-2)/((x+1)(y+1) , y(1)=1

OpenStudy (anonymous):

You need some sort of boundary condition, e.g \(y(0) = \text{something}\)

OpenStudy (anonymous):

\[x ^{2}\frac{ dy }{dx }=\frac{ 4x ^{2}-x-2 }{ (x+1)(y+1)},y(1)=1\]

OpenStudy (anonymous):

okay so i've integrated both sides and no i'm kind of stuck

OpenStudy (anonymous):

1/2 y^2+y=3ln(x+1)+lnx+2/x+c

OpenStudy (anonymous):

This is a non-linear equation. Are you sure you've written it right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its the equation i posted above

OpenStudy (anonymous):

tears as i stare at the screen

OpenStudy (anonymous):

Ok, you've done the intergal correctly. You can get \(c\) by plugging in \(y=1\) and \(x=1\) Then using this value of \(c\) you can actually factorise the \(y\) part and proceed t solve.

OpenStudy (anonymous):

did you work this out on paper?

OpenStudy (anonymous):

Iterative typing in expressions into wolfram alpha. I did it step by step though.

OpenStudy (anonymous):

what is the value of c?

OpenStudy (anonymous):

\[-\frac12 - \ln(8)\] if it's on the side with all the \(x\)'s

OpenStudy (anonymous):

3/2=3ln2+ln1+2+c

OpenStudy (anonymous):

\[c = 2-\frac32 - 3\ln(2) -\ln(1)\] \[3\ln(2) = \ln(2^3) = \ln(8)\] \[\ln(1) = 0\]

OpenStudy (anonymous):

oops, I meant \(\frac32 - 2\) at the start there

OpenStudy (anonymous):

so c=-1/2-ln8, do we just leave it as it is?

OpenStudy (anonymous):

\[\frac2x+\ln(x)+3\ln(1+x) - \ln(8) = \frac12 (y^2 + 2y + 1)\] This should now be solveable

OpenStudy (anonymous):

im sorry where did the +1 come from on the right sight and why did you multiply by 1/2

OpenStudy (anonymous):

Ok, so we got as far as \[\frac2x+\ln(x)+3\ln(x+1)-\ln(8)-\frac12 = \frac12 y^2 + y\] you can move the \(-\frac12\) over to the other side \[\frac2x+\ln(x)+3\ln(x+1)-\ln(8) = \frac12 y^2 + y + \frac12\] factorise out the \(\frac12\) \[\frac2x+\ln(x)+3\ln(x+1)-\ln(8) = \frac12 (y^2 + 2y+ 1)\] one step further: you can factorise the expression in \(y\) \[\frac2x+\ln(x)+3\ln(x+1)-\ln(8) = \frac12 (y+1)^2\]

OpenStudy (anonymous):

bless YOUR SOUL!

OpenStudy (anonymous):

could you possibly finish the problem?

OpenStudy (anonymous):

my brain is going to explode

OpenStudy (anonymous):

but I COMPLETELY UNDERSTOOD what you just wrote

OpenStudy (anonymous):

I'll tell you what to do: multiply both sides by 2 square root both sides subtract 1 from both sides and you're done

OpenStudy (anonymous):

and then plug in 1 for both x and y?

OpenStudy (anonymous):

You've already done that bit to get \(c\)

OpenStudy (anonymous):

i can SLEEP at night now!